2 mole equimolar mixture of Na2C2O4 and H2C2O4 required V1 litre of 0.1 M KMnO4 in acidic medium for complete oxidation . The same amount of mixture required V2 litre of 0.2 M NaOH for neuterlization. The ratio of V1:V2 is

Grade:11

3 Answers

Vikas TU
7 years ago
From the equivalants concept,
moles x n-factor = constant.
Therfore,
(1*2) + (1*2) = 0.1*V1*5
we get,
V1 = 8
For V2,
2 + 2 = 0.2*V2*1
V2 = 4/0.2 => 20
Ratio V1/V2 = 8/20 => 2/5
or
2:5
PSK
4 years ago
Answer will not be 2:5 as NaOH won't react with Na2C2O4 as it is a salt but will only react with H2C2O4.
Ravinder Gahlawat
2 years ago
Answer will be 4:5 as NaOH only reacts with H2C2O4
 
Therefore, 1*2 = 0.2 * v2 * 1
 
             V2 = 10 
 
 
 
 
 
Jxjxjjcjcjjxjxjnsnxjjcjxjjsjsjcnjcjjxjsjjsmdnjcjcjxjjsjsmxncjcjc